OpenStudy (anonymous):

Step 1: |x-2|+3=7 Step 2: |x-2|=7-3 Step 3: |x-2|=4 Which of the following is a correct next step to solve the equation? A) x+2=-4 B) -x-2=4 C) x+2=4 D) x-2=-4 Please help this makes no sense!!

OpenStudy (solomonzelman):

after |x-2|=4 there are two options for x. either x-2=4 or x-2=-4

OpenStudy (solomonzelman):

There isnt a single answer.

OpenStudy (mathstudent55):

Therefore, the only choice that includes one of the two options is the answer.